# DRESS: A Continuous Framework for Structural Graph Refinement

### v1

@@ 1 Introduction
The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation. As we show in Section 3, Original-DRESS distinguishes graphs that 1-WL cannot. However, it still fails to distinguish highly symmetric structures such as Strongly Regular Graphs (SRGs).
@@
Throughout this paper, the graph fingerprint is the sorted vector of converged edge values $\operatorname{sort}(d^{*})$. Two graphs are declared non-isomorphic if and only if their fingerprints differ element-wise beyond a numerical tolerance $\epsilon$. This applies uniformly to all DRESS variants.
@@ 2 Original-DRESS
The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation:
@@
Self-loops. Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$). The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbours would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
Initialization. The equation is scale-invariant (degree-0 homogeneous), so the fixed point is independent of the initial scale. Any uniform initialisation $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$. In practice, $c=1$ is used.
However, because Original-DRESS aggregates strictly over triangles (common neighbours), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
@@ 3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
DRESS operates on edges and aggregates over common neighbours (triangles). We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
$K_{3,3}$: No two adjacent vertices share a common neighbour ($K_{3,3}$ is triangle-free). Hence every edge has the same neighborhood structure, and by symmetry the unique fixed point assigns the same value to all 9 edges. Numerically, $d^{*}_{K_{3,3}}\approx 1.155$ for all edges.
Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbour) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbours besides self-loops). Suppose for contradiction that all 9 edges converge to the same value $d^{*}$. Then:
@@ 4 Motif-DRESS
Original-DRESS is limited to triangle neighborhoods. What if we use other structural motifs? Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional weight function $w(e)$:
@@
The weights $w(e)$ act as a multiplicative factor, controlling how much structural information flows along each edge. Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
@@ 4.1 Properties
Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists. For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot\Delta)$. For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
@@ 4.3 Proof Sketch of Convergence
A complete formal verification of the contraction constant is deferred to future work; all empirical tests confirm convergence within 20 iterations. ∎
@@ 6.2 Properties
Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot\Delta)$, where $I$ is the number of iterations per deletion. The computation for each deleted subgraph is entirely independent, making $\Delta$-DRESS embarrassingly parallel.

### v2

# arXiv:2602.20833v2[cs.DS] 26 Feb 2026

@@ 1 Introduction
The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbor aggregation. As we show in Section 3, Original-DRESS distinguishes graphs that 1-WL cannot. However, it still fails to distinguish highly symmetric structures such as Strongly Regular Graphs (SRGs).
@@
Throughout this paper, the graph fingerprint is the sorted vector of converged edge values $\operatorname{sort}(d^{*})$. Two graphs are declared non-isomorphic if and only if their fingerprints differ element-wise beyond a numerical tolerance $\epsilon$. This applies uniformly to all DRESS variants. An equivalent representation is the histogram $h(d^{*})$: since all DRESS values lie in $[0,2]$ and the convergence tolerance is $\epsilon=10^{-6}$, each value maps to one of $\lceil 2/\epsilon\rceil=2\times 10^{6}$ integer bins, yielding a fixed-size bin-count vector that uniquely identifies the same multiset. In general, the number of bins is determined by the convergence tolerance: a tolerance of $\epsilon$ over the range $[0,2]$ gives $\lceil 2/\epsilon\rceil$ bins. The histogram is particularly useful when fingerprints from multiple subgraphs must be pooled (see $\Delta$-DRESS, Section 6).
@@ 2 Original-DRESS
The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbor aggregation.
@@
Self-loops. Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$). The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbors would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
Initialization. The equation is scale-invariant (degree-0 homogeneous), so the fixed point is independent of the initial scale. Any uniform initialization $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$. In practice, $c=1$ is used.
However, because Original-DRESS aggregates strictly over triangles (common neighbors), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
@@ 3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
DRESS operates on edges and aggregates over common neighbors (triangles). We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
$K_{3,3}$: No two adjacent vertices share a common neighbor ($K_{3,3}$ is triangle-free). Hence every edge has the same neighborhood structure, and by symmetry the unique fixed point assigns the same value to all 9 edges. Numerically, $d^{*}_{K_{3,3}}\approx 1.155$ for all edges.
Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbor) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbors besides self-loops). Suppose for contradiction that all 9 edges converge to the same value $d^{*}$. Then:
@@ 4 Motif-DRESS
Original-DRESS is limited to triangle neighborhoods. What if we use other structural motifs? Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional symmetric weight function $w:E\to\mathbb{R}_{>0}$:
@@
The symmetric weight function $w(e)$ acts as a multiplicative factor, controlling how much structural information flows along each edge. Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
@@ 4.1 Properties
Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists. For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot T)$, where $T$ is the number of triangles. For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
@@ 4.3 Proof Sketch of Convergence
A complete formal verification of the contraction constant is deferred to future work; all empirical tests in this paper confirm convergence within 20 iterations. ∎
@@ 6.2 Properties
Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot T)$, where $T$ is the number of triangles and $I$ is the number of iterations per deletion. The computation for each deleted subgraph is entirely independent, making $\Delta$-DRESS embarrassingly parallel.